OpenStudy (anonymous):

solve x^2y^5 over xy^6

OpenStudy (anonymous):

\[\large \frac{x^2y^5}{xy^6}\] cancel away

OpenStudy (anonymous):

\[\large \frac{x^{\cancel{2}^1}y^5}{\cancel{x}y^6}\]

OpenStudy (anonymous):

\[=\frac{xy^5}{y^6}=\frac{xy^{\cancel{5}}}{y^{\cancel{6}^1}}=\frac{x}{y}\]